OpenStudy (anonymous):

The perimeter of a rectangular cardboard is 500cm and its area is 14400cm^2 . Find the dimensions of the cardboard.

Parth (parthkohli):

\(\color{Black}{\Rightarrow 2(a +b) = 500 }\) \(\color{Black}{\Rightarrow ab = 14400 }\) Solve.

OpenStudy (anonymous):

How to solve ? o-o ? Help please

OpenStudy (anonymous):

Haha ParthKohli already made it simple for you. Oh well. 2(a+b) = 500 a+b = 250 -(1) ab= 14400 -(2) Take eqn. (1), and rearrange it to, a= 250 - b Substitute the above into eqn.2 (250-b) b = 14400 250b - b^2 = 14400 b^2 -250b + 14400 = 0 Solve for b. Once you get b then substitute back to (1) and find a then. VOila answer.

OpenStudy (anonymous):

If you want to get b the only way is to use the general equation of quadratic formula??

OpenStudy (anonymous):

Or factorization?

OpenStudy (anonymous):

Factorization will work, but as you are working in very big numbers it is tricky to find the right ones. Both will work fine.

OpenStudy (anonymous):

Okays thanks alot :)